OpenStudy (anonymous):

Which pair is equivalent? A) 6/10 and 9/25 B) 3/8 and 15/35 C) 14/21 and 12/18 D) 2/5 and 5/20

OpenStudy (anonymous):

c) since 14/21 = 2/3 12/18 = 2/3